### v1

# Random Long-Context Access for Mamba via Hardware-aligned Hierarchical Sparse Attention

Abstract
A key advantage of Recurrent Neural Networks (RNNs) over Transformers is their linear computational and space complexity enables faster training and inference for long sequences. However, RNNs are fundamentally unable to randomly access historical context, and simply integrating attention mechanisms may undermine their efficiency advantages. To overcome this limitation, we propose Hierarchical Sparse Attention (HSA), a novel attention mechanism that enhances RNNs with long-range random access flexibility while preserving their merits in efficiency and length generalization. HSA divides inputs into chunks, selecting the top- chunks and hierarchically aggregates information. The core innovation lies in learning token-to-chunk relevance based on fine-grained token-level information inside each chunk. This approach enhances the precision of chunk selection across both in-domain and out-of-domain context lengths. To make HSA efficient, we further introduce a hardware-aligned kernel design. By combining HSA with Mamba, we introduce RAMba, which achieves perfect accuracy in passkey retrieval across 64 million contexts despite pre-training on only 4K-length contexts, and significant improvements on various downstream tasks, with nearly constant memory footprint. These results show RAMba’s huge potential in long-context modeling. 00footnotetext: ‡ Corresponding authors
1 Introduction
The success of Large Language Models (LLMs) [11, 1, 56] has been largely driven by the Transformer architecture [57]. However, the quadratic computational and memory costs of self-attention make it inefficient for processing long sequences. Moreover, Transformers often struggle with inputs that exceed their pre-training length. These limitations have renewed interest in alternative architectures such as Recurrent Neural Networks (RNNs) [10, 24, 32, 27] that enable efficient, linear-time processing of sequential data while retaining a degree of extrapolation capability.
However, RNN-based models suffer from a critical limitation: an information bottleneck [55] caused by compressing variable-length contexts into fixed-dimensional representations. Unlike attention mechanisms, they lack random access to contextual information, which becomes especially problematic in tasks like passkey retrieval, where performance degrades as sequence length increases [59]. While augmenting RNNs with attention mechanisms can help mitigate this limitation, it introduces new challenges—such as poor length extrapolation, quadratic computational complexity, and substantial memory overhead during inference—ultimately undermining the original efficiency advantages of RNNs. Consequently, there remains no satisfactory RNN-based solution that can simultaneously achieve length generalization, random-access flexibility, and efficiency.
To address the trilemma, we propose a novel Hierarchical Sparse Attention (HSA) mechanism. Existing sparse attention methods [66, 31] typically divide sequences into chunks, allowing each token to attend to a concatenation of selected chunks, achieving efficiency while retraining random access capability. While this design offers a promising step toward resolving the trilemma, a closer examination reveals a critical weakness: these methods often suffer from inaccurate chunk selection, both within the training distribution and when generalizing to longer, out-of-domain contexts. As illustrated in Figure 1(A), the issue may stem from estimating token-to-chunk relevance via attention scores, which are learned from token-to-token gradients instead of feedback from entire chunks.
Based on this insight, HSA introduces a two-stage hierarchical mechanism to enable end-to-end learning of token-to-chunk relevance. As shown in Figure 1(B), in the first stage, it separately weights over tokens within each chunk to capture chunk-level information. In the second stage, it fuses the chunk information with weights derived from token-to-chunk scores. During the backpropagation pass, the chunks that are more helpful for next-token prediction are assigned higher weights. Experimental results show that this chunk selection mechanism remains accurate even when context lengths exceed pre-training lengths by 1000 times under a certain condition. Since each token corresponds to different chunks, a naive implementation would require substantial memory. Thus, we further propose a hardware-aligned algorithm to achieve efficient parallel computation.
Built upon HSA, we propose Random Access Mamba (RAMba), an extension of Mamba-2 [16] that incorporates HSA into specific layers. To maintain a constant memory footprint during inference, RAMba offloads the token-level key-value (KV) cache to CPU memory while retaining compact chunk-level representations on the GPU for efficient chunk selection. At each time step, only the KV cache of the selected chunks is loaded onto the GPU, ensuring efficient memory usage. To further minimize GPU-CPU memory swaps, RAMba leverages the hidden states derived from an intermediate layer as shared KV cache for all subsequent HSA layers, requiring just one chunk selection and swap per step. This architecture can simultaneously balance training and inference efficiency, length generalization ability, and long-range random access flexibility.
In our experiments, we compare RAMba against baselines like Transformers, Mamba-2, their variants with sliding window attention [14] and NSA [66] across long-range language modeling, downstream tasks, and efficiency. RAMba consistently outperforms the baselines in long-context modeling and downstream tasks while exhibiting exceptional length generalization. Notably, it is the first Mamba-based model to achieve perfect accuracy on a 64M context in the passkey retrieval task. In terms of efficiency, HSA is three times faster than NSA and 5–25 times faster than flash attention with 16K context length or longer in forward pass. Additionally, With memory offloading enabled, RAMba maintains nearly constant memory usage. These results demonstrate RAMba’s superior capability in long-text modeling. In summary, our contributions are three fold:
1.
We propose HSA, a novel hierarchical attention mechanism with a hardware-efficient algorithm, offering efficiency, length generalization, and long-range random access flexibility.
2.
Based on HSA, we introduce RAMba, which integrates the advantages of the attention mechanism into Mamba while maintaining a nearly constant memory footprint during inference.
3.
We conducted comprehensive experiments on the length generalization of Mamba with various attention mechanisms. The results show that HSA excels in both performance and efficiency.
2 Related Works
Sparse Attentions.
Sparse attention aims to reduce computational complexity by focusing on a limited number of tokens. For example, sliding window attention and its variants [40, 14, 67, 9] restrict computations to a fixed-size local window for each token. Such methods often sacrifice the ability to capture long-range information. Clustering based approaches [29, 58, 61, 51] employs locality-sensitive hashing or K-Means for token clustering and perform attention with clusters. These methods often trade off efficiency for quality due to the limited accuracy of clustering. Cache eviction approaches [70, 21, 30] maintain constant memory costs by retaining only the most important tokens in the KV cache, but this limits the model’s ability to access arbitrary contexts randomly. Combiner [49] utilizes a hierarchical attention mechanism by compressing tokens within chunks into single key-value pairs via max-pooling. uowever, this approach sacrifice the ability to capture token-level information inside the chunk, resulting in reduced accuracy. Recently, NSA [66] and MoBA [31] achieve sparsity by dividing a sequence into chunks and dynamically selecting relevant chunks for each token based on summed token-level attention scores. However, these approaches struggle with accurately identifying important chunks. Our work introduces a two-stage hierarchical attention mechanism with sparse chunk selection, achieving end-to-end learning of token-to-chunk relevance, while maintaining the ability to capture token-level information, resulting in attention flexibility, accuracy, and efficiency.
RNN-base Language Models.
RNN-based language models have recently gained increased attention because their per-token inference cost remains constant as the sequence length grows. Linear Attention [27] shows that replacing softmax attention with kernel-based approximations allows Transformers to be reformulated as RNNs, achieving similar performance while benefiting from recurrent properties. Recent advancements in RNNs, such as RWKV series [41, 42], Mamba series [23, 17], GLA [63], HGRN series [45, 46], xLSTM [8] and others [69, 64], continue to push the boundaries of this approach. Our work builds on these RNN models and is in principle applicable across different RNN architectures. Meanwhile, other studies [18, 36, 50] explore hybrid architectures of RNNs and attention by leveraging for long-range memory and attention for short-range patterns. In contrast, our approach leverages RNNs to adaptively retain essential short-range information for next-token prediction, while sparse attention selectively accesses long-range context at arbitrary positions.
Context Length Extrapolation.
The full attention mechanism struggles to generalize to contexts longer than those observed during pretraining [44], even with techniques like relative positional encoding [52] or attention weight scaling [60]. While methods such as Landmark Attention [34] and DCA [3] have been proposed to address this limitation, they still encounter a sharp rise in perplexity when extrapolating beyond a certain length, typically 32. Recently, GCA [26] achieves perfect accuracy on 16M context lengths despite being pre-trained on only 4K lengths by integrating end-to-end retrieval within attention. However, GCA limits retrieval to once every (usually 64) tokens, reducing its flexibility. The two-stage attention mechanism in HSA is inspired by GCA, but achieving per-token retrieval through kernel optimization.
3 Methodology
From the psychological perspective, human memory primarily comprises two main systems: working memory [6], which temporarily stores and manipulates information with limited capacity, and long-term memory [5], which stores information indefinitely with virtually unlimited capacity. Inspired by this, RAMba applies Mamba to simulate working memory by compressing variable-length contexts into a finite representation for manipulation. Meanwhile, it uses HSA to model long-term memory as an extendable KV cache, enabling efficient retrieval and attention computation. In HSA, the key innovation lies in its two-stage hierarchical attention mechanism: first, weighting over tokens within a chunk, and then weighting over chunks. By offloading KV cache to CPU memory or disk, it can theoretically maintain unlimited memory. In the following sections, we elaborate the model architecture, HSA, kernel design, and optimizations for training and inference.
3.1 Model Architecture
As shown in Figure 2(a), RAMba contains Mamba layers, equally divided into upper and lower decoders. Given the input sequence , their corresponding embeddings are fed into the lower decoder and the outputs are divided into chunks according to chunk size . These chunked representations are then passed through a Transformer-based bi-directional encoder independently to form chunked memories, used for chunk selection and sparse attention in subsequent HSA layers. The upper decoder takes the outputs of the lower decoder as inputs, in which HSA layers alternate with Mamba layers, repeating times. HSA layers capture long-range dependencies by attending to the selected chunks. Between the lower and upper decoders, a chunk selection layer selects the most relevant chunks based on the dot-product similarity between token and chunk representations.
Notations.
We focus on architectures like GQA [2], where every query heads shares a common KV head. The hidden size is divided into groups, satisfying , with . Here, is the dimension of each group, and is the dimension per head. Since all groups perform chunk selection and HSA identically, subsequent notation refers to an arbitrary group representative of all. Specifically, denotes one group of the output of the -th layer at step , and denotes one group of the encoded token representations of the -th chunk.
Chunk Selection.
At each time step , each token uses the representation derived from the lower decoder to select top- past chunks. Each query group independently selects these chunks, which are shared across the heads within the group. Formally, for each group, we have:
where is RMSNorm [68], denotes the ranking position in descending order, is the mean-pooled representations within the -th chunk, and represents token and chunk representations for chunk selections, is the relevance score of to the -th chunk, and is the selected chunk indices for .
3.1.1 Hierarchical Sparse Attention
Chunk Weights.
To avoid the impact of position encoding on length generalization [28], we discard it and instead model the ordinal relationships of distances akin to the stick-breaking attention [53]. In the stick-breaking process, participants sequentially take a portion from a stick, with later participants dividing the remaining portion left by earlier ones, thus introducing a “most recent” bias. Analogously, let the total weights serve as the stick and denote the weight assigned to the -th chunk selected by the -th token, we have:
where is the proportion taken from the remaining attention, calculated via the sigmoid function . The function rearranges in descending order to prioritize chunks closer to in attention allocation. denotes the the -th value in .
Hierarchical Attention.
After obtaining the weights for each chunk, performs attention with tokens in each retrieved chunk separately. We denote the representation of information collected from the -th selected chunk as . These representations are then fused using the chunk weights. Formally, for the -th HSA layer, we have:
where are the key and value corresponding to the -th chunk, which are shared across all HSA layers. is the query representation of at the -th layer and is obtained by applying RMSNorm [68] followed by a linear transformation.
3.1.2 Kernel Design
In HSA, each token corresponds to a distinct group of chunks, which can lead to a substantial memory footprint in a naive implementation. Inspired by NSA, we address this issue by implementing hardware-aligned HSA kernels based on Triton [54]. As illustrated in Figure 2(b), each GPU thread loads the query representations for a single step along with the selected chunks’ KV pairs for attention computation. The Algorithm 1 demonstrates a parallel forward pass for one group of query heads, with multiple groups processed in parallel following the same approach. Particularly, we use softmax-off-by-one [33] to allow tokens in the current chunk to ignore any retrieved tokens. During the back-propagation process, we adopt a two-phased backward pass inspired by Dao [15]. The first stage accumulates gradients for and , followed by and in the second stage. The pseudo-code for this process is shown in Algorithm 2 and 3, in which denotes whether the -th token selectes the -th chunk, and represents the index of the -th chunk in
3.2 Training & Inference
Training.
Many efforts have been made to address Mamba’s length generalization issues. A simple yet effective solution is truncated backpropagation through time (BPTT) [13, 63], in which the first state of a sequence is initialized as the final state of its preceding sequence. We follow this approach in training RAMba and other baselines. However, even though RAMba demonstrates certain extrapolation capabilities, we still observe an increase in perplexity on longer contexts. We hypothesize that RNNs’ memory state provides certain shortcuts [22] for long-range attention, degrading performance on contexts largely exceeding the pre-trained length. Thus we attempt to introduce appropriate forgetting mechanism into the memory state to disrupt the shortcuts. A straightforward way is memory reset, where sequences are divided into equal segments, and the initial state of each segment is reset to zero. To align with BPTT, we set the initial state as the last hidden state of a random segment in the previous step. Our experiments show that this straightforward technique effectively improves length generalization for RAMba.
Inference.
A key challenge in introducing attention mechanisms to RNNs is managing the memory footprint, as the KV cache scales linearly with sequence length. Prior works [34, 26] have demonstrated the feasibility of offloading the KV cache to CPU memory and selectively loading chunks during inference. For a prompt of length , all chunk representations are offloaded to the CPU after prefilling, while only is kept in GPU memory for chunk selection. During decoding, RAMba retrieves and loads chunks at each step. Since the KV cache is shared across all HSA layers, the number of parameters exchanged between the CPU and GPU totals , with per group head number usually set to 16. Our efficiency analysis in § 4.4 demonstrates the overhead for memory exchange is fully acceptable. Since the memory footprint of still increase with the sequence length, to theoretically achieving constant memory, a straight forward way is to offloading it to a FAISS [19] database. However, in practice, the memory footprint of is very limited, making such offloading unnecessary. Detailed analysis is elaborated in the experiments section.
4 Experiments
To ensure a fair comparison, we pre-train all models from scratch with 4K context length to observe their performance and extrapolation capabilities across various tasks.
4.1 Setups
Baselines.
We adopt the Mamba-2 architecture as the backbone of the RNN model and YaRN [43] as the Transformer baseline. The parameter size of all models trained from scratch is 370M, with detailed parameters provided in Appendix A. We study Mamba variants with different attention mechanisms, including sliding window attention, native sparse attention (NSA), and HSA. For sliding window attention, the window size is set to 512, incorporating two position encoding schemes: ALibi [44] and RoPE, the latter following the settings in Samba [50]. We set the chunk size of HSA to 64 following NSA. To ensure that the field of view for sparse attention matches the sliding window size (64 * 8 = 512), we set the number of selected chunks to 8. Regarding the NSA, we use its efficient open-source implementation 111https://github.com/fla-org/native-sparse-attention. To purely compare the capabilities of the sparse attention components, we disable NSA’s sliding window attention. The HSA incorporates a single-layered Transformer-based bi-directional encoder for chunk memory encoding, accounting for 5.4% of the total parameters. As it does not participate in decoding, its impact on fairness is minimal. HSA layers are inserted into the upper decoder every Mamba layers, while the placement of other attention layers remains consistent with this pattern. These settings remain consistent across all subsequent 370M models. Since the compressed attention in NSA functions similarly to Combiner [49], we don’t conduct a separate comparison against it. Some other related works [35] are not included in the experiments due to the lack of a well-implemented open-source reproduction.
Memory Reset.
We use w/ m.r. and w/o m.r. to denote whether the memory reset is enabled during training. Specifically, for an 8K token context, the memory is reset every 4K tokens, which aligns with the context length of other baselines. However, the chunk selection scope for sparse attention spans 8K tokens, which might be unfair to other baselines. To ensure fair comparisons, we apply the same settings to both HSA and NSA.
Pre-training.
4.2 Long-range Langauge Modeling
Datasets.
We evaluate long-range language modeling on the recognized datasets in the field, which include PG19 [47], ArXiv-math [4], and CodeParrot [62].
| Models(370M) | pg19 | arxiv | code | pg19 | arxiv | code | pg19 | arxiv | code |
| eval_len=4k | eval_len=16k | eval_len=64k | |||||||
| Transformer | 18.61 | 4.23 | 3.28 | 539.15 | 199.42 | 62.17 | >104 | >104 | 2865.51 |
| Mamba | 17.92 | 4.24 | 3.28 | 17.38 | 3.91 | 3.09 | 17.30 | 3.86 | 3.05 |
| w/ SWA | 17.82 | 4.21 | 3.26 | 20.48 | 5.01 | 3.53 | 23.86 | 6.46 | 3.96 |
| w/ SWA | 17.82 | 4.21 | 3.26 | 17.50 | 4.03 | 3.19 | 17.80 | 4.25 | 3.35 |
| w/ NSA | 17.87 | 4.20 | 3.25 | 17.31 | 3.87 | 3.06 | 17.31 | 3.87 | 3.05 |
| w/ NSA | 17.74 | 4.18 | 3.24 | 17.56 | 4.29 | 3.26 | 17.62 | 4.35 | 3.28 |
| RAMba | 17.82 | 4.15 | 3.23 | 17.15 | 3.73 | 3.04 | 17.01 | 3.65 | 3.07 |
| RAMba | 17.63 | 4.13 | 3.21 | 17.11 | 3.81 | 3.08 | 17.11 | 3.87 | 3.21 |
Results.
From Table 1, RAMba performs better than the baselines across most datasets, both in-domain and out-domain. In addition, we have two findings. First, Mamba’s perplexity decreases as the context length increases from 4K to 64K, whereas Mamba with attentions show higher perplexity at 64K compared to 16K when not trained with memroy reset. This suggests that Mamba’s memory state may influence the generalization ability of attention mechanisms. Secondly, when trained with memory reset, Mamba with NSA or HSA all exhibit stronger length extrapolation capabilities but show a decline in in-domain performance. This result validates the effectiveness of memory reset, which essentially constrains the model to rely entirely on the content of the text for retrieval, thereby preventing the model from learning shortcuts. Since the model becomes adapted to retrieving information from longer contexts, it is reasonable that its performance declines when handling shorter contexts.
4.3 Downstream Tasks
Tasks.
We evaluate various models’ long-context modeling abilities on classic tasks like passkey retrieval [34] and the LongBench V2 dataset [7]. To increase task difficulty, we replace numbers in passkey retrieval with random token sequences. Since passkey retrieval is relatively simple, we further fine-tuned the models using synthetic data following RULER [25]. We use a context length of 4K for fine-tuning with a total training step size equivalent to 5% of the pre-training stage. Evaluations were conducted across different lengths on four RULER tasks: Single NIAH (S-N), Multi-queries NIAH (MQ-N), Variable Tracking (VT), and Frequent Words Extraction (FWE). To align with passkey retrieval, keys in Single NIAH were also replaced with random token sequences. We follow Waleffe et al. [59] by using a Cloze format for LongBench evaluation. As LongBench V2 is a zero-shot benchmark thus small models may exhibit randomness, we additionally evaluate on fine-tunable datasets, including summarization like XSUM [38] and CNN [37], and QA tasks like SQuaD [48], HotpotQA [65], and QuALITY [39]. NSA’s implementation currently does not support generative tasks, so its generative results were not reported.
| Models(370M) | Overall | Easy | Hard | Short | Medium | Long |
| Transformer | 24.8 | 23.4 | 25.7 | 28.5 | 19.6 | 29.5 |
| Mamba | 22.4 | 16.4 | 26.1 | 19.6 | 22.7 | 26.3 |
| w/ SW | 23.3 | 26.9 | 21.0 | 19.6 | 26.8 | 22.1 |
| w/ SW | 24.2 | 25.7 | 23.2 | 20.3 | 26.3 | 26.3 |
| w/ NSA | 23.5 | 25.7 | 22.1 | 21.5 | 25.8 | 22.1 |
| w/ NSA | 22.1 | 26.3 | 19.6 | 17.1 | 26.3 | 22.1 |
| RAMba | 25.7 | 28.7 | 23.9 | 20.9 | 27.8 | 29.5 |
| RAMba | 23.5 | 26.3 | 21.7 | 20.9 | 23.7 | 27.4 |
Results.
First, Figure 3 shows that RAMba achieves perfect accuracy in the passkey retrieval task with a 64M context, even without memory reset, while most baselines drop to nearly zero around 64K. However, removing the chunk encoder causes RAMba’s accuracy to decrease as context length grows, highlighting the encoder’s importance for HSA and length generalization. One possible reason is that the hidden states are primarily optimized for predicting the next token, thus requiring an adapter to extract information representing the current context. Second, LongBench’s evaluation results in Table 2 are generally consistent with those of language modeling. However, since LongBench is a zero-shot evaluation dataset, most baselines performed below the random guessing rate of 25%, with the results across different sub-tasks show high variability. Thus, we believe following SFT tasks can provide deeper insights into architectural capabilities for small models.
| Models (370M) | S-N | MQ-N | VT | FWE | S-N | MQ-N | VT | FWE | S-N | MQ-N | VT | FWE | S-N | FWE |
| ctx-len=4K | ctx-len=64K | ctx-len=256K | ctx-len=1M | |||||||||||
| Transformer | 95.08 | 88.59 | 97.12 | 60.02 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | — | — |
| Mamba-2 | 96.66 | 1.86 | 61.78 | 64.75 | 10.45 | 0.00 | 8.96 | 37.31 | 0.00 | 0.00 | 6.25 | 43.75 | 0.00 | 40.00 |
| w/ SW | 91.84 | 6.96 | 37.20 | 72.08 | 28.36 | 0.00 | 17.91 | 0.00 | 6.25 | 0.00 | 0.00 | 0.00 | — | — |
| w/ NSA (w/o m.r.) | 92.58 | 59.09 | 57.70 | 66.98 | 13.43 | 0.00 | 0.00 | 2.99 | 6.25 | 0.00 | 0.00 | 6.25 | — | — |
| w/ NSA (w/ m.r.) | 84.79 | 4.08 | 31.63 | 49.54 | 4.48 | 0.00 | 5.97 | 26.87 | 0.00 | 0.00 | 0.00 | 12.50 | — | — |
| RAMba | 91.74 | 80.61 | 95.18 | 49.81 | 85.07 | 55.22 | 55.22 | 53.73 | 62.50 | 62.50 | 37.50 | 37.50 | 24.00 | 20.00 |
| RAMba | 92.76 | 87.10 | 96.66 | 76.81 | 55.22 | 29.85 | 68.66 | 8.96 | 18.75 | 12.50 | 31.25 | 0.00 | — | — |
| RAMba | 92.67 | 81.08 | 57.79 | 75.97 | 41.79 | 23.88 | 14.92 | 26.87 | 6.25 | 12.50 | 0.00 | 6.25 | — | — |
Table 3 presents the evaluation results of various models fine-tuned on RULER synthetic data across different context lengths. We have four interesting findings. First, HSA demonstrates more precise chunk selection. HSA performs comparably to full-attention on retrieval-related tasks (Single-NIAH, MultiQuery-NIAH, Variable Tracking) over in-domain context length, while NSA lags behind both, supporting our argument that estimating chunk importance using token-to-token attention scores is inaccurate. Meanwhile, HSA achieves strong performance across retrieval tasks despite selecting chunks only once, validating the effectiveness of the hierarchical attention mechanism in learning token-to-chunk relevance. Second, Mamba excels in sequential statistic tasks. While Mamba-based models underperform in retrieval tasks, they outperform in Frequent Word Extraction (FWE), successfully extrapolating up to 256 the pretraining length. This advantage likely stems from Mamba’s continuous memory flow mechanism, which is also partially inherited by RAMba. Third, memory reset helps length generalization. Models with memory reset consistently outperform those without in length extrapolation. Though performance drops significantly beyond 1M context length on more challenging retrieval tasks, still achieving 256 extrapolation. Fourth, Single-NIAH extrapolates only to 4M, while the similar task of passkey retrieval extends to 64M, with the sole difference lying in the prompt pattern as passkey retrieval uses a simpler prompt 222The passkey retrieval template is structured as follows: “(essays) The pass key is ¡PASS KEY¿. (essays) What is the passkey? The passkey is”. The single NIAH template follows this format: “(essays)… One of the special magic numbers for long-context is: ¡PASS KEY¿. (essays)… What is the special magic number for long-context mentioned in the provided text? Answer:” which may facilitate learning more generalizable patterns, underscoring the potential of HSA. Still, precise chunk selection for extremely long contexts remains an open challenge for future work.
| Models (370M) | XSUM | CNN | SQuaD | HotpotQA | QuALITY | AVG. |
| avg-len=498 | avg-len=883 | avg-len=174 | avg-len=1428 | avg-len=7745 | ||
| Mamba-2 | 30.40/11.89/24.53 | 37.97/17.54/35.84 | 41.33/52.03 | 18.70/26.20 | 33.32 | 31.03 |
| w/ SWA | 30.78/12.11/24.79 | 38.96/17.84/36.73 | 50.54/61.23 | 19.96/27.58 | 29.48 | 32.57 |
| w/ SWA | 30.88/12.33/24.99 | 38.78/17.67/36.57 | 51.46/61.91 | 19.89/28.00 | 32.89 | 32.61 |
| w/ NSA | — | — | — | — | 31.40 | — |
| Transformer | 30.10/11.63/24.23 | 38.25/17.89/36.14 | 45.50/56.13 | 21.90/29.49 | 33.46 | 32.82 |
| RAMba | 30.81/12.25/24.80 | 39.11/18.04/36.85 | 48.24/59.17 | 22.30/30.53 | 34.13 | 33.64 |
Table 4 shows the results of the model after fine-tuning on summarization and generation tasks. RAMba still maintains a lead in the vast majority of tasks, but it are two observations worth noting. First, Mamba with SWA demonstrates an advantage over other models on the SQuaD task. One potential reason is that the SQuaD dataset contains shorter texts even less than the sliding window, making SWA able to randomly access all contexts. This phenomenon suggests that even short-range random access poses certain challenges for RNNs. Second, even though Transformer underperforms Mamba in PPL, its downstream task performance remains strong, second only to RAMba. This further validates the importance of random access capability for downstream tasks.
4.4 Efficiency Analysis
Experimental Setup.
To evaluate the scalability and efficiency of HSA, we evaluate the runtime and memory footprint of Flash-Attention, NSA, and HSA operators across different sequence lengths, the throughput of various models at different scales, and the per token time consumption during inference. Runtime measurements were conducted with three attention layers only, excluding additional components such as MLPs or Mamba. HSA performed a single chunk selection shared across all three layers, while the sliding window of NSA is still disabled. The memory footprint was reported as a multiple of the Mamba 370M model. We use w/ offloading and w/o offloading to denote whether memory offloading is enabled. To further evaluate the benefits of shared chunk selection, we introduce an ablation group (w/o sharing), where chunk selection and CPU-GPU memory happens in each HSA layer. When measuring training throughput, we enable FSDP [71] and gradient checkpointing [12], running models on 16 Physics Processing Units (PPUs), each with approximately half the computational power of an A100 GPU.
| Models | Parameter Size | |||
| 370M | 780M | 1.4B | 3B | |
| Mamba | 16.44 | 9.71 | 6.53 | 3.46 |
| w/ SWA | 15.64 | 9.24 | 6.27 | 3.31 |
| w/ full_attn | 10.00 | 5.73 | 4.31 | 2.07 |
| w/ NSA | 14.35 | 8.70 | 5.46 | 3.18 |
| RAMba | 14.80 | 8.76 | 5.70 | 3.18 |
| Models | Prompt-Length | ||
| 4K | 16K | 64K | |
| Transformer | 2.26 | 8.90 | 32.12 |
| Mamba-2 | 2.92 | 2.82 | 2.84 |
| RAMba | 3.14 | 3.05 | 2.76 |
| RAMba | 3.97 | 4.19 | 4.02 |
| RAMba | 5.98 | 5.87 | 5.95 |
Results.
Figure 4 compares the time consumption and memory footprint of three operators across different context lengths. Both NSA and HSA, as sparse attention mechanisms, significantly outperform Flash-Attention in terms of speed. HSA is faster than NSA because it performs the chunk selection only once, its only operation with quadratic complexity, and shares it across all HSA layers. In contrast, NSA involves computations with quadratic complexity like compressed token attention and chunk selection in every layer, which increases its time consumption. For memory footprint, enabling memory offloading drastically reduces GPU memory usage and slows its growth with increasing context length. Despite introducing CPU-GPU memory exchange, the impact on inference speed is limited, as reported in Table 6. For example, with a chunk size of 64, hidden size of 4096 and batch size of 16, the total offloaded data is , which is acceptable given CPU-GPU bandwidth exceeding 20GB/s. Table 6 shows that RAMba achieves 90% of Mamba’s training throughput. Although HSA is faster than NSA, the additional encoder introduced sometimes offsets this advantage at certain scales. Nonetheless, the results clearly highlight HSA’s high efficiency, excellent scalability during training, and near-constant memory usage during inference.
5 Conclusion
In this work, we propose a novel hierarchical sparse attention (HSA), which demonstrates excellent training and inference efficiency, while also features advantages such as length extrapolation and long-ragne random access flexibility, making it an ideal architecture for long-text modeling. It achieves perfect accuracy in passkey retrieval with a 64M context, showcasing its potential for extrapolating to infinite lengths. However, in RULER experiments with more complex tasks, accuracy still declines with increasing length, though it significantly outperforms baseline models. How to ensure precise chunk selection on complex tasks over infinite lengths remains a challenge for future research.
Appendix A Hyper-parameters
| Architecture | Transformer | Mamba | Mamba+NSA | Mamba+SWA | RAMba |
| Total Params (M) | 372 | 368 | 375 | 375 | 385 |
| Hidden size, | 1024 | 1024 | 1024 | 1024 | 1024 |
| Mamba Layers | - | 48 | 48 | 48 | 48 |
| Attention Layers | 16 | - | 3NSA | 3SWA | 3HSA |
| Other Layers | 16MLP | - | - | - | 1Chunk Selection, 1Encoder |
| MLP hidden size | 5504 | - | - | - | 1344 |
| Query heads | 32 | - | 16 | - | 16 |
| KV heads | 16 | - | 1 | - | 1 |
| Vocab size | 50280 | 50280 | 50280 | 50280 | 50280 |
Appendix B Training hyper-parameters
All 370M models used the AdamW optimizer with
linear learning rate warmup with warmup ratio , cosine decay to .
peak learning rate .
total tokens , batch size tokens.
gradient clip value 1.0
no dropout
no linear bias terms
weight decay
AdamW hyperparameter (the GPT3 value)